If class mark of a class interval is 8.5the class size is 5 then the class limit of the corresponding class interval is

Grade:9

1 Answers

BALA SRIVATSAV YERRAMILLI
3 years ago

Heyy!

Given: Class mark ‘x’ = 8.5;
           Class size ‘h’ = 5.
 
Therefore, the lower limit is:
L = x -  h/2 = 8.5 – 2.5 = 6.
 
And the upper limit is:
U = x +h/2 = 8.5 +2.5 = 11
Therefore, the class interval = 6 - 11.

Hope this helps you! :)